OpenStudy (anonymous):

identify the vertex and the axis of symmetry of the graph of the function y = 3(x + 2)2 – 3. (1 point)vertex: (2, –3); axis of symmetry: x = 2 vertex: (–2, –3); axis of symmetry: x = –2 vertex: (2, 3); axis of symmetry: x = 2 vertex: (–2, 3); axis of symmetry: x = –2

OpenStudy (anonymous):

@ghazi

OpenStudy (ghazi):

have you tried to plot it?

OpenStudy (anonymous):

idk what it even is ghazi

OpenStudy (ghazi):

y=3x+6-3=3x+3 is this clear?

OpenStudy (anonymous):

combine like terms?

OpenStudy (ghazi):

yes, now lets say we have to find value of y at different values of X

OpenStudy (anonymous):

well y = 3x+3=3x-+3

OpenStudy (ghazi):

y=3x+3, now when x=0,y=3 and when x=-1, y=0 we use this to plot a graph

OpenStudy (ghazi):

wait what is your function it looks like parabola?

OpenStudy (anonymous):

huh?

OpenStudy (ghazi):

\[y=(3x+2)^2-3\] is this your function?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

no wait the 3 is outside the ()

OpenStudy (ghazi):

\[y=3(x+2)^2-3\] this one

OpenStudy (anonymous):

yes